Suppose the interest rate on a 1-year T-bond is 5.0% and that on a 2-year T-bond is 7.0%. Assuming the pure expectations theory is correct, what is the market's forecast for 1-year rates 1 year from now?

a. 7.36%
b. 7.75%
c. 8.16%
d. 8.59%
e. 9.04%
Answer #1
Let Face Value of bond = 1
value of 1-year bond , reinvested after one year , at the end of year 2
= value of 2-year bond at the end of year 2

=> 1*(1+0.05)(1+k) = 1*(1+0.07)^2 (where, k = rate of 1-year bond at year 1)
=> k = 0.0938 = 9.38% = 9.04% (approx.)

Hence, ANSWER is OPTION (e)